OpenStudy (anonymous):

Find the degree of the polynomial functions or if they are not a polynomial function. 1. 5x^10 + 13x^4 -17x^3 2. 5x^3y^2 + 13x^4y^3-17x^3y^4 3. x^7 + 3 4. x^4 + 3x^2 -5x 5. 5x^2 - 3x^5 -5x^3 Answer Choices 7 9 10 5 4 Not a polynomial in 1 variable.

OpenStudy (anonymous):

Are you matching the choices to the numbers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ah ok well the degree of a polynomial is the exponent to the very leading part of the function so like: 3x^2+x-8, 2 would be the degree. Does that help?

OpenStudy (anonymous):

a little, it's still confusing to me though

OpenStudy (anonymous):

ok ill do the first for u. so we have 5x^10 + 13x^4 -17x^3 correct? Well that would match the choice of 10 since the leading portion is 5x^10 and the exponent is 10 the degree of it is 10. Help?

OpenStudy (anonymous):

Ok, so the second would be 4 and the third would be 7?

OpenStudy (anonymous):

the third one would be 7 yes but the second one i am not so sure about i didn't deal with functions involving more then 1 variable and picking out a degree

OpenStudy (anonymous):

so the second would not be a polynomial in 1 variable, right?

OpenStudy (anonymous):

would the fourth be 4 and the fifth be 5?

OpenStudy (anonymous):

now that i rethink this the 2nd is 7 because you add up the exponents of each term so it would be: 2+3=5, 4+3=7, 3+4=7 and since 7 is the highest degree for a term it is the highest for the function. The third one i believe is the not a polynomial in 1 variable. The 4th and 5th you got correct

OpenStudy (anonymous):

Ah I get it thank you

OpenStudy (anonymous):

Welcome!